OpenStudy (anonymous):

What is the electric power used by an appliance if the current is 2 A and the voltage is 50 volts? Show work and include all units

OpenStudy (anonymous):

We will assume it is a DC powered appliance. The equation for electrical power is\[P = V \cdot I\]where V is the voltage and I is the current.

OpenStudy (anonymous):

P=V*I P=50*2 P=100 watts

OpenStudy (anonymous):

For your edification, that power formula comes from the relationship between potential energy and voltage. Let us move an infinitesimal charge \(dq\) through a voltage drop \(V\), which releases a certain amount of energy \(dW\).\[dW=V\ dq \Rightarrow \frac{dW}{dt}=V\ \frac{dq}{dt} \Rightarrow P=VI\]

OpenStudy (anonymous):

the formula for power is \[P =V I\] wherein V is for potential difference (SI unit volt) and I is current flowing through the conductor (SI unit Ampere).thus answer for your question is:- \[P =V I\] P=2A*50V P=100Watts.

OpenStudy (radar):

It did not specify DC so AC should be considered also. There was good reason to avoid AC and hope it was just DC. If it was AC we would need to know the phase lag between voltage and current (power factor)

OpenStudy (anonymous):

@radar Wouldn't averaging over a period still turn out the same?

OpenStudy (radar):

Phase difference between voltage and current due to reactive components takes place in the appliance. Then simply multiplying VI produces what is known as apparent power. This apparent power consist of real power that is consumed by the load, and a "circulating" power that travels back and forth between load and source. The phase angle termed theta is called the "power factor" and is equal to the cosine of the angle theta so the Real power would be P=kVI where k=cosin theta. The power factor is desired to be 1 or no phase angle and both energy producers and users strive to minimize the angle. It should be pointed out that the imaginary portion of apparent power is bad as it does result in losses in the transmitting wires.